OpenStudy (anonymous):

i need help on simplifying the fallowing expression 7(a-9) and (4-z)15 all steps plz

OpenStudy (mathteacher1729):

It looks like you must distribute. 7(q + 2) = 7*q + 7*2 = 7q + 14 or 8(g-3) = 8*g - 8*3 = 8g -24

OpenStudy (anonymous):

this is the rong equation i meant 7 (a-9) and (4-z) 15

OpenStudy (anonymous):

he has tried to explain how it should be done you can follow his directions for your question n solve it yourself